OpenStudy (anonymous):

What is the equation of the line that has a slope of 3 and goes through point (4, 7)?

OpenStudy (anonymous):

y=3x+b 7=3(4)+b 7=12+b b=-5 y=3x-5

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

np